# arXiv:2603.03156v1[math.NT] 03 Mar 2026

### v1

@@ Abstract
This is a survey article on the twisted dynamical zeta functions of Ruelle and Selberg and the Fried’s conjecture. It is based on the mini-course: "Twisted Ruelle zeta function, complex-valued analytic torsion and the Fried’s conjecture", given by the author during the thematic trimester programme: "Representation Theory and Noncommutative Geometry" at Institut Henri Poincaré.
@@ 1.2 Fried’s conjecture
Are there special vaues for $R(s)$? Secondly, one may want to consider representations of the lattice $\Gamma$, since holonomy representations are often considered. Moreover, these dynamical zeta functions can be studied under different and more complicated geometrical considerations, for example, higher-dimensional hyperbolic manifolds with cusps, orbifolds, general locally symmetric spaces and Riemannian manifolds of variable negative curvature.
@@
In the case of an arbitrary representation of the fundamental group of a compact hyperbolic manifold, the resolution of the Fried’s conjeture 1.2.1 was given by the author, under certain assumptions ([52]) by Frahm and the author ([17]) and by Frahm, Bénard and the author([4]). More precisely, we have the following theorems.
@@ 1.2.3 Odd dimensional compact hyperbolic manifolds
([50, Theorem 7.13]) For every $\sigma\in\widehat{M}$, the Ruelle zeta function $R(s;\sigma,\chi)$ admits a meromorphic continuation to the whole complex plane $\mathbb{C}$.
@@
([52, Theorem 1]) Let $\chi$ be a finite-dimensional complex representation of $\Gamma$. Let $\Delta^{\sharp}_{\chi,k}$ be the flat Hodge Laplacian, acting on the space of $k$-differential forms on $X$ with values in the flat vector bundle $E_{\chi}$. Then, the Ruelle zeta function has the representation
@@
Let $\operatorname{Rep}(\pi_{1}(X),\mathbb{C}^{n})$ be the set of all $n$-dimensional, complex representations of $\pi_{1}(X)$. This set has a natural structure of a complex algebraic variety. Let $V\subset\operatorname{Rep}(\pi_{1}(X),\mathbb{C}^{n})$ be an open neighbourhood (in classical topology) of the set $\operatorname{Rep}^{u}_{0}(\pi_{1}(X),\mathbb{C}^{n})$ of acyclic, unitary representations such that, for all $\chi\in V$, the odd signature operator $B_{\chi}$ is bijective (see [52, Section 3] and [6]).
@@
([52, Theorem 2]) Let $\chi\in V$. Then, the Ruelle zeta function $R(s;\chi)$ is regular at $s=0$ and is equal to the complex Cappell-Miller torsion,
@@ 1.2.4 Related work
Bunke-Olbrich in [7] studied the twisted dynamical zeta functions associated with unitary representations for all locally symmetric spaces of real rank one and proved the Fried’s conjecture. In the higher rank case and for unitary representations the Fried’s conjecture was proved by Shen ([43]), see also Moscovici-Stanton ([31]), Moscovici-Stanton-Frahm ([32]). For other contributions to Fried’s conjecture and also applications of its resolution, see Müller ([34], [30], [35]), Pfaff ([37]), Shen ([44], [45], [47], [46]), Shen-Yu ([47]), Bismut-Shen ([5]), Deitmar ([13]), Hochs-Saratchandran ([24]), Hochs-Pirie ([25]), Dyatlov-Zworski ([15]), Guillarmou-Dang-Rivière, Shen ([11]), Chaubet-Dang ([10]) and Cekić-Delarue-Dyatlov-Paternain ([9]), Jorgenson-Lee-Smajlovic ([27]), Malo-Zung ([26]) and Humbert-Tao ([53]). We mention also here a related work by Abdurrahman-Venkatesh ([1]).
@@ 2.3.1 Real hyperbolic manifolds of odd dimension
We consider compact real hyperbolic manifolds $X$ of odd dimension $d$. By [50, Lemma 3.3], we have the following lemma.
@@
Using (2.2) and (2.3), we can define the twisted dynamical zeta functions for non-unitary representations $\chi$ (for more details see [50, Section 3]).
@@ 2.3.2 Compact hyperbolic surfaces
Let now $X$ be compact hyperbolic surface. To define the twisted dynamical zeta functions for non-unitary representations $\chi$, we use the fact that the word length of an element in $\Gamma$ is quasi isometric to the hyperbolic distance (see [29, Proposition 3.2] and [56, p.10]). We have that there exists a $c\geq 0$ such that
@@
    The constants $c_{1}$ and $c_{2}$ can be also determined explicitly in terms of the term $|\rho|$, and the constant $\beta$ in (2.2). In fact, Lemma 2.3.1 is also a consequence of the result of Lubotzky, Mozes and Raghunathan ([29, Proposition 3.2]).
@@ 3 Trace formulas
For the derivation of the trace formula, we follow the work of Wallach ([54]), and its extension to non-unitary representations setting by Müller ([33]). We denote by $\operatorname{spec}(\Delta^{\sharp}_{\chi})\subseteq\mathbb{C}$ the discrete) spectrum of $\Delta^{\sharp}_{\chi}$. For $\mu\in\operatorname{spec}(\Delta^{\sharp}_{\chi})$, we write $L^{2}(X,E_{\chi})_{\mu}$ for the corresponding generalized eigenspace. We define the algebraic multiplicity $m(\mu)$ of $\mu$ as $m(\mu):=\dim L^{2}(X,E_{\chi})_{\mu}$. We use the extended results for the heat operator obtained in [50, p. 171–173] to conclude that the heat operator $e^{-t{\Delta}^{\sharp}_{\chi}}$ is an integral operator with smooth kernel, i.e., there exists a smooth section $H^{\chi}_{t}$ of $\operatorname{End}(E_{\chi})$ on $X\times X$ such that for $f\in L^{2}(X,E_{\chi})$, we have
@@
By [33, Proposition 2.5], this operator is of trace class. By Lidskii’s theorem [49, Theorem 3.7], we have
@@
where $x=\Gamma gK,x^{\prime}=\Gamma g^{\prime}K\in X$ with $g,g^{\prime}\in G$. By [50, Proposition 5.3], we have the following result.
@@
As in [33, Proposition 5.1, Proposition 6.1] and [55, p. 172–173, 177–178], we group the summation into conjugacy classes and use the Fourier inversion formula to obtain
@@ 3.0.2 Compact hyperbolic odd-dimensional manifolds
(Trace formula, [50, Theorem 5.5]) Let $\sigma\in\widehat{M}$.
@@ 4.2 Special values
where the sum is taken over representatives $\gamma_{j}$ of the $r$ inconjugate elliptic classes with orders $\nu_{j}$, $\chi(X)$ is the Euler characteristic of $X$, and $\zeta$ denotes the Riemann zeta function (see [28, Section 6]).
@@
Moreover, by [28, Corollary 8.1.1], we have that if $\mu=0$ is not the eigenvalue of the twisted Laplacian, then the twisted Selberg zeta function $Z(s;\rho)$ is non-vanishing at $s=1$ and
@@
In [28] also, the authors studied the asymptotic behaviour of the torsion factor for a sequence of non-unitary representations introduced by Yamaguchi in [57] and prove that the asymptotic behaviour of this constant as the dimension of the representation tends to infinity is the same as the behaviour of the higher-dimensional Reidemeister torsion on the unit tangent bundle (up to an absolute constant). In particular, by [28, Corollary 7.2], we have

### v2

# Twisted dynamical zeta functions and the Fried’s conjecture

@@ Abstract
This is a survey article on the twisted dynamical zeta functions of Ruelle and Selberg and the Fried’s conjecture. It is based on the mini-course: "Twisted Ruelle zeta function, complex-valued analytic torsion and the Fried’s conjecture", given by the author during the thematic trimester programme: "Representation Theory and Noncommutative Geometry" at the Institut Henri Poincaré.
@@ 1.2 Fried’s conjecture
Are there special values for $R(s)$? Secondly, one may want to consider representations of the lattice $\Gamma$, since holonomy representations are often considered. Moreover, these dynamical zeta functions can be studied under different and more complicated geometrical considerations, for example, higher-dimensional hyperbolic manifolds with cusps, orbifolds, general locally symmetric spaces and Riemannian manifolds of variable negative curvature.
@@
In the case of an arbitrary representation of the fundamental group of a compact hyperbolic manifold, the resolution of the Fried’s conjecture 1.2.1 was given by the author, under certain assumptions ([53]) by Frahm and the author ([17]) and by Frahm, Bénard and the author([4]). More precisely, we have the following theorems.
@@ 1.2.3 Odd dimensional compact hyperbolic manifolds
([51, Theorem 7.13]) For every $\sigma\in\widehat{M}$, the Ruelle zeta function $R(s;\sigma,\chi)$ admits a meromorphic continuation to the whole complex plane $\mathbb{C}$.
@@
([53, Theorem 1]) Let $\chi$ be a finite-dimensional complex representation of $\Gamma$. Let $\Delta^{\sharp}_{\chi,k}$ be the flat Hodge Laplacian, acting on the space of $k$-differential forms on $X$ with values in the flat vector bundle $E_{\chi}$. Then, the Ruelle zeta function has the representation
@@
Let $\operatorname{Rep}(\pi_{1}(X),\mathbb{C}^{n})$ be the set of all $n$-dimensional, complex representations of $\pi_{1}(X)$. This set has a natural structure of a complex algebraic variety. Let $V\subset\operatorname{Rep}(\pi_{1}(X),\mathbb{C}^{n})$ be an open neighbourhood (in classical topology) of the set $\operatorname{Rep}^{u}_{0}(\pi_{1}(X),\mathbb{C}^{n})$ of acyclic, unitary representations such that, for all $\chi\in V$, the odd signature operator $B_{\chi}$ is bijective (see [53, Section 3] and [6]).
@@
([53, Theorem 2]) Let $\chi\in V$. Then, the Ruelle zeta function $R(s;\chi)$ is regular at $s=0$ and is equal to the complex Cappell-Miller torsion,
@@ 1.2.4 Related work
Bunke-Olbrich in [7] studied the twisted dynamical zeta functions associated with unitary representations for all locally symmetric spaces of real rank one and proved the Fried’s conjecture. In the higher rank case and for unitary representations the Fried’s conjecture was proved by Shen ([45]), see also Moscovici-Stanton ([33]), Moscovici-Stanton-Frahm ([32]). For other contributions to Fried’s conjecture and also applications of its resolution, see Müller ([35], [31], [36]), Pfaff ([38]), Shen ([46], [47], [48]), Shen-Yu ([44]), Bismut-Shen ([5]), Deitmar ([13]), Hochs-Saratchandran ([25]), Hochs-Pirie ([24]), Dyatlov-Zworski ([15]), Guillarmou-Dang-Rivière, Shen ([11]), Chaubet-Dang ([10]) and Cekić-Delarue-Dyatlov-Paternain ([9]), Jorgenson-Lee-Smajlovic ([28]), Malo-Zung ([27]) and Humbert-Tao ([26]). We mention also here a related work by Abdurrahman-Venkatesh ([1]).
@@ 2.3.1 Real hyperbolic manifolds of odd dimension
We consider compact real hyperbolic manifolds $X$ of odd dimension $d$. By [51, Lemma 3.3], we have the following lemma.
@@
Using (2.2) and (2.3), we can define the twisted dynamical zeta functions for non-unitary representations $\chi$ (for more details see [51, Section 3]).
@@ 2.3.2 Compact hyperbolic surfaces
Let now $X$ be compact hyperbolic surface. To define the twisted dynamical zeta functions for non-unitary representations $\chi$, we use the fact that the word length of an element in $\Gamma$ is quasi isometric to the hyperbolic distance (see [30, Proposition 3.2] and [56, p.10]). We have that there exists a $c\geq 0$ such that
@@
    The constants $c_{1}$ and $c_{2}$ can be also determined explicitly in terms of the term $|\rho|$, and the constant $\beta$ in (2.2). In fact, Lemma 2.3.1 is also a consequence of the result of Lubotzky, Mozes and Raghunathan ([30, Proposition 3.2]).
@@ 3 Trace formulas
For the derivation of the trace formula, we follow the work of Wallach ([54]), and its extension to non-unitary representations setting by Müller ([34]). We denote by $\operatorname{spec}(\Delta^{\sharp}_{\chi})\subseteq\mathbb{C}$ the discrete) spectrum of $\Delta^{\sharp}_{\chi}$. For $\mu\in\operatorname{spec}(\Delta^{\sharp}_{\chi})$, we write $L^{2}(X,E_{\chi})_{\mu}$ for the corresponding generalized eigenspace. We define the algebraic multiplicity $m(\mu)$ of $\mu$ as $m(\mu):=\dim L^{2}(X,E_{\chi})_{\mu}$. We use the extended results for the heat operator obtained in [51, p. 171–173] to conclude that the heat operator $e^{-t{\Delta}^{\sharp}_{\chi}}$ is an integral operator with smooth kernel, i.e., there exists a smooth section $H^{\chi}_{t}$ of $\operatorname{End}(E_{\chi})$ on $X\times X$ such that for $f\in L^{2}(X,E_{\chi})$, we have
@@
By [34, Proposition 2.5], this operator is of trace class. By Lidskii’s theorem [50, Theorem 3.7], we have
@@
where $x=\Gamma gK,x^{\prime}=\Gamma g^{\prime}K\in X$ with $g,g^{\prime}\in G$. By [51, Proposition 5.3], we have the following result.
@@
As in [34, Proposition 5.1, Proposition 6.1] and [55, p. 172–173, 177–178], we group the summation into conjugacy classes and use the Fourier inversion formula to obtain
@@ 3.0.2 Compact hyperbolic odd-dimensional manifolds
(Trace formula, [51, Theorem 5.5]) Let $\sigma\in\widehat{M}$.
@@ 4.2 Special values
where the sum is taken over representatives $\gamma_{j}$ of the $r$ inconjugate elliptic classes with orders $\nu_{j}$, $\chi(X)$ is the Euler characteristic of $X$, and $\zeta$ denotes the Riemann zeta function (see [29, Section 6]).
@@
Moreover, by [29, Corollary 8.1.1], we have that if $\mu=0$ is not the eigenvalue of the twisted Laplacian, then the twisted Selberg zeta function $Z(s;\rho)$ is non-vanishing at $s=1$ and
@@
In [29] also, the authors studied the asymptotic behaviour of the torsion factor for a sequence of non-unitary representations introduced by Yamaguchi in [57] and prove that the asymptotic behaviour of this constant as the dimension of the representation tends to infinity is the same as the behaviour of the higher-dimensional Reidemeister torsion on the unit tangent bundle (up to an absolute constant). In particular, by [29, Corollary 7.2], we have